# Kerr-Newman black hole in a Melvin-swirling universe

### v1

@@ 1 Introduction
After the construction of the Schwarzschild and Kerr-swirling solutions [5], some other extensions have been discovered [7, 8, 9, 10, 11, 12, 13, 14], and many physical properties of the swirling universe have been studied [15, 16, 17, 18, 19, 20], proving that these spacetimes are a fertile subject of investigation.
@@
The main tool for embedding black holes in the swirling universe is the Ehlers transformation [21]. As it is well known, the Einstein-Maxwell equations for a stationary and axisymmetric metric can be rewritten in an equivalent form, the celebrated Ernst equations [22, 23]. The latter make manifest some Lie point symmetries, which allow us to map a solution of the Einstein-Maxwell system into another one. The Ehlers map is one of these symmetries, and its action in the context of the swirling universe was clarified in [5] Here, we apply the Ehlers transformation to the dyonic Kerr-Newman-Melvin metric [6] to add the swirling rotation, thus obtaining a general solution that encompasses many interesting sub-cases.
The structure of the present paper is as follows. In Sec. 2 we present the Ernst equations and the Ehlers transformation needed to construct the novel solution. In Sec. 3, we present the dyonic Kerr-Newman-Melvin-swirling spacetime and investigate its properties. In particular, we study the singularity structure of the solution, focusing on the tuning of the parameters which allows us to completely remove the conical singularities and the Dirac strings. Such a choice, as we show in Sec. 3.1.1, removes the typical curvature singularity hidden behind the horizon, thus leaving us with a completely regular manifold. We also list, for the various sub-cases, when it is possible to choose the parameters to remove the singularities and when it is not. Thus, the general solution that we construct offers the possibility to systematize the swirling family and to clarify the role of the parameters.
@@ 3 Dyonic Kerr-Newman in a Melvin-swirling universe
The spacetime (3.1) represents an electromagnetically charged and rotating black hole, embedded in a swirling and Melvin background. As such, it is characterized by six parameters: the mass $M$, the charges $Q$ and $H$, the angular momentum $a$, the external magnetic field $B$ and the swirling parameter $\jmath$.
@@ 3.1 Singularities
In particular, we will show that the new spacetime (3.1) is completely regular outside the event horizon according to an appropriate choice of the parameters: a constraint on the parameters will allow us to remove both the conical singularities and the Dirac string (which naturally appear in the presence of a magnetic charge). In addition, this choice of parameters will also make the new spacetime free of curvature singularities, while CTCs will still be present, but hidden behind the event horizon.
@@ 3.1.2 Conical singularities
To compute the angular defects, we consider the ratio between the perimeter of a small circle around the symmetry axis and its radius [5], for both halves of the symmetry axis, $\theta=0$ and $\theta=\pi$
@@ 3.1.3 Dirac strings
For the other sub-cases, we have that if both the Melvin and swirling parameters are present, the Dirac strings are always removable or never present, except if the black hole is rotating but not charged, which is quite a surprising result since the Dirac strings usually arise in black holes that do carry a magnetic charge. On the other hand, if only the swirling parameter is present, it is possible to have Dirac strings if any type of charge is present, which however is an expected result since, in a certain sense, the addition of the swirling rotation generates a magnetic charge from an electric one, and vice-versa [^3] In Appendix is reported the dyonic Reissner-Nordström black hole in a swirling universe with a clarification on this topic..
@@ 3.1.4 Closed timelike curves
with functions $F$ and $\Sigma$ respectively defined by Eq. (3.2) and Eq. (3.8).
It can be proved numerically that $F\geq 0$ for any choice of the parameters if $M,r\geq 0$. Thus, we have that the regions in which CTCs are allowed are determined by the condition
$\Sigma=\bigl{(}r^{2}+a^{2}\bigr{)}^{2}-\bigl{(}r^{2}-2Mr+Q^{2}+H^{2}+a^{2}\bigr{)}\,a^{2}\sin^{2}\theta<0\,,$ (3.28)
which does not depend on either the Melvin parameter $B$ or the swirling parameter $\jmath$. This implies that these regions are the same as in the case without the Melvin-swirling background.
@@ 3.1.5 Misner strings
Misner strings [28, 29], which are the gravitational analog of the Dirac strings, appear in rotating spacetimes, when the non-diagonal function $\omega=g_{t\phi}/g_{tt}$ is not regular on the symmetry axis, taking different values on the two hemispheres $\theta=0$ and $\theta=\pi$. Furthermore, the presence of a Misner string also leads to CTCs and causality issues.
@@ 3.2 Curvature
Finally, we study the curvature of by analyzing the Kretschmann scalar $K=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. For large radial distances, $r\to\infty$, the Kretschmann scalar of the dyonic Kerr-Newman metric  decays as
@@ 4 Ergoregions and geometry of the horizons
We will now analyze the ergoregions, i.e. the regions of spacetime where the frame dragging makes it impossible for an observer to remain static, and the event horizons. Moreover, we will also provide various plots, in order to perform a qualitative analysis of these quantities. More specifically, we will start by analyzing the dyonic Kerr-Newman black hole in a Melvin-swirling universe (3.1), with the singularities removed through Eqs. (3.24).
@@
where $g_{(1)}$ and $g_{(1)}$ are some differentiable functions, and where we recall that $\Delta=0$ defines the locations of the horizons (3.15). From this result, we can conclude that the poles of the horizons ($\Delta=0$, $\cos^{2}\theta=1$) can never be part of the ergoregion, since in these points $g_{tt}\neq 0$. Moreover, it can be proved that on the symmetry axis ($\cos^{2}\theta=1$) it holds $g_{(2)}<0$. Therefore, we have that the symmetry axis is never part of the ergoregions outside the outer horizon and also inside the inner horizon, since in these regions $\Delta>0$. Analogously, the portion of the symmetry axis between the outer and inner horizons is always part of the ergoregions, since between the horizons we have $\Delta<0$.
Furthermore, it holds that $g_{(1)}=0$ if $a=B=0$. Thus, we can conclude that $g_{tt}=0$ for the entirety of the horizons ($\Delta=0$) for all the non-rotating black holes in a swirling universe sub-cases, independently from the specific value of the parameters. Due to this last consideration, we will also present plots for the dyonic Reissner-Nordström black hole in a swirling universe in Appendix F, using the same values of the parameters, except for the magnetic field $B$, which is fixed by the constraints needed to remove the conical singularities for the Melvin-swirling black hole (C.5), while it is obviously zero for the swirling sub-case.
@@ 5 Near-horizon extremal geometry and Kerr/CFT correspondence
$\hat{A}$ $=\frac{\hat{A}_{0}}{r_{0}^{2}}\,\hat{r}\,d\hat{t}+\frac{2\pi}{\delta\phi}\frac{\hat{A}_{3}}{\Gamma}\Bigl{(}d\hat{\phi}+\tilde{\kappa}\,\hat{r}\,d\hat{t}\Bigr{)}\,,$ (5.4b)
@@
$\tilde{\kappa}$ $=-\frac{\delta\phi}{2\pi}\frac{\kappa}{r_{0}^{2}}\,,\qquad\gamma(\theta)=\frac{2\pi}{\delta\phi}\frac{r_{0}^{2}\sin\theta}{\Gamma}\,,\qquad\alpha(\theta)=1\,,$ (5.5b)
@@
and thus it equals the Bekenstein-Hawking entropy as computed in (3.58). We stress that the computation works only if the constraints (3.24) hold: this means that the the above quantities must be evaluated for the regular solution [^6] The various quantities are not written explicitly in terms of the regularised parameters for sake of brevity..
@@ Appendix A Dyonic Kerr-Newman in a Melvin universe
The above functions defined as the expansion of the following quantities:

### v2

@@ 1 Introduction
After the construction of the Schwarzschild and Kerr-swirling solutions [5], some other extensions have been discovered [7, 8, 9, 10, 11, 12, 13], and many physical properties of the swirling universe have been studied [14, 15, 16, 17, 18, 19], proving that these spacetimes are a fertile subject of investigation.
@@
The main tool for embedding black holes in the swirling universe is the Ehlers transformation [20]. As it is well known, the Einstein-Maxwell equations for a stationary and axisymmetric metric can be rewritten in an equivalent form, the celebrated Ernst equations [21, 22]. The latter make manifest some Lie point symmetries, which allow us to map a solution of the Einstein-Maxwell system into another one. The Ehlers map is one of these symmetries, and its action in the context of the swirling universe was clarified in [5]. Here, we apply the Ehlers transformation to the dyonic Kerr-Newman-Melvin metric [6] to add the swirling rotation, thus obtaining a general solution that encompasses many interesting sub-cases.
The structure of the present paper is as follows. In Sec. 2, we present the Ernst equations and the Ehlers transformation needed to construct the novel solution. In Sec. 3, we present the dyonic Kerr-Newman-Melvin-swirling spacetime and investigate its properties. In particular, we study the singularity structure of the solution, focusing on the tuning of the parameters which allows us to completely remove the conical singularities and the Dirac strings. Such a choice, as we show in Sec. 3.1.1, removes the typical curvature singularity hidden behind the horizon, thus leaving us with a completely regular manifold. We also list, for the various sub-cases, when it is possible to choose the parameters to remove the singularities and when it is not. Thus, the general solution that we construct offers the possibility to systematize the swirling family and to clarify the role of the parameters.
@@ 2 Generation of the solution
It is interesting to note that, in this construction, the Ehlers and Harrison transformations commute. In fact, it is possible to obtain our result by applying the Harrison map to the swirling Kerr-Newman solution. This is a highly non-trivial result, since the Ehlers and Harrison maps belong to $\mathrm{SU(2,1)}$, which is not abelian. This point is discussed in [25].
@@ 3 Dyonic Kerr-Newman in a Melvin-swirling universe
The spacetime (3.1) represents an electromagnetically charged and rotating black hole, embedded in a swirling and Melvin background [^1] We remind the reader that the Melvin background represents a spacetime filled with a magnetic field, and it was firstly found and studied in [26, 27].. As such, it is characterized by six parameters: the mass $M$, the charges $Q$ and $H$, the angular momentum $a$, the external magnetic field $B$ and the swirling parameter $\jmath$.
@@ 3.1 Singularities
In particular, we will show that the new spacetime (3.1) is completely regular outside the event horizon according to an appropriate choice of the parameters: a constraint on the parameters will allow us to remove both the conical singularities and the Dirac string (which naturally appear in the presence of a magnetic charge), while the Misner string is not present because there is no NUT parameter. In addition, this choice of parameters will also make the new spacetime free of curvature singularities, while CTCs will still be present, but hidden behind the event horizon.
@@ 3.1.2 Conical singularities
To compute the angular defects, we consider the ratio between the perimeter of a small circle around the symmetry axis and its radius [5], for both halves of the symmetry axis, $\theta=0$ and $\theta=\pi$:
@@ 3.1.3 Dirac strings
For the other sub-cases, we have that if both the Melvin and swirling parameters are present, the Dirac strings are always removable or never present, except if the black hole is rotating but not charged, which is quite a surprising result since the Dirac strings usually arise in black holes that do carry a magnetic charge. On the other hand, if only the swirling parameter is present, it is possible to have Dirac strings if any type of charge is present, which however is an expected result since, in a certain sense, the addition of the swirling rotation generates a magnetic charge from an electric one, and vice-versa [^4] In Appendix B.2 is reported the dyonic Reissner-Nordström black hole in a Melvin-swirling universe with a clarification on this topic in Appendix G..
@@ 3.1.4 Closed timelike curves
with functions $F$ and $\Sigma$ respectively defined by Eq. (3.2) and Eq. (3.8). However, using the argument explained in Appendix H, we have that the regions in which CTCs are allowed do not depend on either the Melvin parameter $B$ or the swirling parameter $\jmath$, and are the same as those of the Kerr-Newman black hole, which are determined only by the condition
$\Sigma=\bigl{(}r^{2}+a^{2}\bigr{)}^{2}-\bigl{(}r^{2}-2Mr+Q^{2}+H^{2}+a^{2}\bigr{)}\,a^{2}\sin^{2}\theta<0\,.$ (3.28)
@@ 3.1.5 Misner strings
Misner strings [30, 31], which are the gravitational analog of the Dirac strings, appear in rotating spacetimes when the non-diagonal function $g_{t\phi}/g_{tt}$ is not regular on the symmetry axis [32], taking different values on the two hemispheres $\theta=0$ and $\theta=\pi$. Furthermore, the presence of a Misner string also leads to CTCs and causality issues.
@@ 3.2 Curvature
Finally, we study the curvature by analyzing the Kretschmann scalar $K=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$. For large radial distances, $r\to\infty$, the Kretschmann scalar of the dyonic Kerr-Newman metric  decays as
@@ 4 Ergoregions and geometry of the horizons
We will now analyze the ergoregions, i.e. the regions for which $g_{tt}>0$, which are the regions of spacetime where the frame dragging makes it impossible for an observer to remain static; and the event horizons. Moreover, we will also provide various plots, in order to perform a qualitative analysis of these quantities. More specifically, we will start by analyzing the dyonic Kerr-Newman black hole in a Melvin-swirling universe (3.1), with the singularities removed through Eqs. (3.24).
@@
where $g_{(1)}$ and $g_{(1)}$ are some differentiable functions, and where we recall that $\Delta=0$ defines the locations of the horizons (3.15). From this result, we can conclude that the poles of the horizons ($\Delta=0$, $\cos^{2}\theta=1$) can never be part of the ergoregion, since in these points $g_{tt}=0$. Moreover, it can be proved that on the symmetry axis ($\cos^{2}\theta=1$) it holds $g_{(2)}<0$. Therefore, we have that the symmetry axis is never part of the ergoregions outside the outer horizon and also inside the inner horizon, since in these regions $\Delta>0$. Analogously, the portion of the symmetry axis between the outer and inner horizons is always part of the ergoregions, since between the horizons we have $\Delta<0$.
Furthermore, it holds that $g_{(1)}=0$ if $a=B=0$. Thus, we can conclude that $g_{tt}=0$ for the entirety of the horizons ($\Delta=0$) for all the non-rotating black holes in a swirling universe sub-cases, independently from the specific value of the parameters. Due to this last consideration, we will also present plots for the dyonic Reissner-Nordström black hole in a swirling and Melvin-swirling universe in Appendix F, using the same values of the parameters, except for the magnetic field $B$, which is fixed by the constraints needed to remove the conical singularities for the Melvin-swirling black hole (C.5), while it is obviously zero for the swirling sub-case.
@@ 5 Near-horizon extremal geometry and Kerr/CFT correspondence
$\hat{A}$ $=\frac{\hat{A}_{0}}{r_{0}^{2}}\,\hat{r}\,d\hat{t}+\frac{2\pi}{\delta\phi}\frac{\hat{A}_{3}}{\Gamma(\theta)}\Bigl{(}d\hat{\phi}+\tilde{\kappa}\,\hat{r}\,d\hat{t}\Bigr{)}\,,$ (5.4b)
@@
$\tilde{\kappa}$ $=-\frac{\delta\phi}{2\pi}\frac{\kappa}{r_{0}^{2}}\,,\qquad\gamma(\theta)=\frac{2\pi}{\delta\phi}\frac{r_{0}^{2}\sin\theta}{\Gamma(\theta)}\,,\qquad\alpha(\theta)=1\,,$ (5.5b)
@@
and thus it equals the Bekenstein-Hawking entropy as computed in (3.58). We stress that the computation works only if the constraints (3.24) hold [^7] Metric (5.4) in the presence of conical singularities may be diffeomorphic to the near-horizon metric of the accelerating Kerr-Newman black hole, according to the conjecture stated in [13]. If this is the case, such a diffeomorphism might be used to compute the mass and the angular momentum of our solution. We leave this point, however, for future improvements.: this means that the the above quantities must be evaluated for the regular solution [^8] The various quantities are not written explicitly in terms of the regularised parameters for sake of brevity..
@@ Appendix A Dyonic Kerr-Newman in a Melvin universe
The above functions are defined as the expansion of the following quantities:
@@ Appendix G Near-horizon extremal geometry for the sub-cases
Appendix H Invariance under an Ehlers-Harrison transformation
Using the definition of the Ernst potentials (2.4), we have that the function $f$ in the LWP ansatz is given by
±$\pm f=Re{(\pazocal{E})}+|\Phi|^{2}\,,$ (H.1)
where the positive sign is for the electric ansatz (2.2), while the minus sign is for the magnetic ansatz (2.3).
Consequently, after an Ehlers transformation (2.7), the transformed function $f^{\prime}$ is given by
$f^{\prime}_{\text{Ehlers}}=\frac{f}{|1+i\,\jmath\,\pazocal{E}|^{2}}=\frac{f}{\bigl{(}1-\jmath\,h)^{2}+\jmath^{2}\bigl{(}f^{2}+|\Phi|^{2}\bigr{)}}\,;$ (H.2)
analogously, for a Harrison transformation (2.8), one has
$f^{\prime}_{\text{Harrison}}=\frac{f}{|1-2\alpha^{*}\Phi-|\alpha|^{2}\pazocal{E}|^{2}}\,.$ (H.3)
Therefore, an Ehlers or Harrison transformation, whether magnetic or electric, never changes the sign of $f$ by construction. Thus, after an electric transformation, i.e. when one adds the NUT parameter or electromagnetic charges, the ergoregions are not modified, since $f=-g_{tt}$. Similarly, a magnetic transformation, i.e. when one embeds the seed in a swirling or Melvin background, does not alter the region in which closed timelike curves (CTCs) are allowed, since $f=g_{\phi\phi}$.